3sina=5sinb,. Then find. Tan((a+b)/2)/tan((a-b)/2)=

Grade:11

1 Answers

Kartik
6 years ago
we have,
3sina=5sinb
rearranging it 
(sina/sinb)=(5/3)
now applying componendo and dividendo rule 
we get 
(sina+sinb)/(sina-sinb)=4
now using the identities of trigonometry we get 
(2sin((a+b)/2)cos((a-b)/2)  )/(  2cos((a+b)/2)sin((a-b)/2))=4
this becomes
(tan((a+b)/2)  )/( tan((a-b)/2)=4